OpenStudy (anonymous):

When using the elimination method (addition and subtraction) When do you know when to add or subtract?

OpenStudy (amistre64):

just add at all times ...

OpenStudy (anonymous):

OK So don't bother about subtracting?

OpenStudy (amistre64):

\[5-3~=~5+(-3)\]

OpenStudy (amistre64):

subtraction isnt really an algebraic operation, its just a shortcut notation.

OpenStudy (anonymous):

Thanks! My book is confusing me with the "Use addition and subtraction, for the elimination method" I think they purposelessly make it confusing.

OpenStudy (amistre64):

the idea is simply, to eliminate one of the variables, you need them to be opposite in sign; and then combining like "parts" eliminates that variable

OpenStudy (amistre64):

3x + 2y = 4 x + 5y = -10 if we multiply the 2nd one by -3 we can combine to get rid of the xs 3x + 2y = 4 -3x -15y = 30 --------------- 0x -13y = 34

OpenStudy (amistre64):

once you know a y value ... you dont need to eliminate again. just plug in y=___ into either equation to solve for x

OpenStudy (anonymous):

Then you can solve by dividing 34/13! No subtraction necessary :D I know I was on to something when I never understood subtraction ;P

OpenStudy (amistre64):

:) good luck